OpenStudy (natalie94):

A student scores 71 out of 100 on a test. If the maximum score for the next test is 100 points, what score does the student need to maintain at least a average of 81. The student needs to score at least blank out of 100

OpenStudy (mathstudent55):

If you took 2 tests and have 2 grades, x and y, how do you calculate the average of the grades?

OpenStudy (mathstudent55):

Do you know how to find the average of two numbers?

OpenStudy (faiqraees):

Solve the equation \[\large \rm \frac{71+x}{200}=81\]

OpenStudy (mathstudent55):

@Natalie94 I'm not trying to put you in the spot. The way to help you is for me to find out what you know and for you to ask questions about what you don't. After you post a question, there should be a dialogue between you and the helper. If you do know how to find the average of 2 numbers, say so. If you don't, let me know, and I'll explain it to you.

OpenStudy (mathstudent55):

Of course, every once in a while, someone decides to throw in an answer without any explanation, and just leave the answer posted even if it is incorrect and only creates more confusion.

OpenStudy (faiqraees):

\(\color{#0cbb34}{\text{Originally Posted by}}\) @mathstudent55 Of course, every once in a while, someone decides to throw in an answer without any explanation, and just leave the answer posted even if it is incorrect and only creates more confusion. \(\color{#0cbb34}{\text{End of Quote}}\) Oh Thank you for bring it to my notice.

OpenStudy (faiqraees):

Solve the equation \[\large \rm \frac{71+x}{2}=81\] I hope its rectified now.

OpenStudy (mathstudent55):

I have to first commend you for owning up to your error. Your equation is rectified, but I think the goal of OS is to explain, which is what I am trying to do. You now posted the correct equation, but you offer no explanation. I'm trying to explain how to get to that equation step by step, but the poster does not respond, so I think it's time to just quit for today.

OpenStudy (faiqraees):

\(\color{#0cbb34}{\text{Originally Posted by}}\) @mathstudent55 I have to first commend you for owning up to your error. Your equation is rectified, but I think the goal of OS is to explain, which is what I am trying to do. You now posted the correct equation, but you offer no explanation. I'm trying to explain how to get to that equation step by step, but the poster does not respond, so I think it's time to just quit for today. \(\color{#0cbb34}{\text{End of Quote}}\) That was exactly the reason I posted the equation.

OpenStudy (natalie94):

Sorry about the delay I left my computer open but the answer I got was x=91 . Is that right?

OpenStudy (faiqraees):

Perfect

OpenStudy (natalie94):

Thank you for your help!

OpenStudy (faiqraees):

No problem. But next time, please let the helpers know that you aren't going to be available for a time. That will be highly appreciable.